# Multi-task Photonic Reservoir Computing: Wavelength Division Multiplexing for Parallel Computing with a Silicon Microring Resonator

### v1

@@ 3 Introduction
Over recent years, the development of photonic computing has caught a lot of attention from the scientific community and positioned photonics as a potential technology for novel computing schemes (Huang et al., 2022; McMahon, 2023). Light offers very interesting features that can be exploited for computing processes, such as the capability for massive parallelization. This can be achieved by multiplexing several optical channels using different wavelengths within an available range of the electromagnetic spectrum. This is a well-established technique used for optical communication known as wavelength division multiplexing (WDM). In WDM, a high density of channels can be transmitted through the same medium with relatively low interference between them, ultimately boosting the total communication capacity as several data streams can be transmitted simultaneously. If a similar principle is used for computing, this can be translated into different computing tasks being addressed simultaneously.
@@
These studies on MRR-based TDRC however, focused on investigating the system when addressing a single task at a time. This is where the benefits of WDM can come into play to enhance the computing capacity of photonic RC schemes. For instance, in (Gooskens et al., 2022), the wavelength dimension is used in a photonic RC scheme to enhance the performance of a specific task (nonlinear signal equalization and boolean operations). Similarly, in (Li et al., 2023) WDM is used in an RC based on a Fabry-Perot semiconductor laser to achieve superior throughput and performance in the task of signal equalization in an optical fiber communication link. Lastly, in (Lupo et al., 2023) a photonic RC based on phase modulation and spectral filtering demonstrated the simultaneous operation of two different instances of the same task, which are encoded using frequency combs. In an initial study (Giron Castro et al., 2024c), we combined WDM with MRR-based TDRC to show that this scheme also offers the potential of parallel computing by simultaneously addressing three distinct computing tasks.
In this work, we extend the study in (Giron Castro et al., 2024c) by systematically analyzing the memory and nonlinear capabilities of the proposed WDM MRR-based TDRC scheme and investigating the impact of additional computing tasks in parallel. Finally, we quantify the effect of the phase control in the external waveguide that is used in this scheme as a feedback mechanism. The numerical model of the system is based on the well-studied temporal coupled-mode theory (TCMT).
@@ 4 Multitask WDM MRR-based TDRC
On a high level, the TDRC scheme based on an MRR relies on injecting the input data encoded onto an optical carrier into the MRR at its input port and detecting a non-linearly transformed version at the drop port. An additional feedback connection with an optimized delay is normally included between the through and add port of the MRR to provide additional memory (Donati et al., 2022; Giron Castro et al., 2024b).
@@ 4.1 Setup and frequency allocation
The simulated system is shown in Fig. 1. We consider $M$ wavelength-multiplexed optical channels, each modulated by the masked input sequence corresponding to a particular task. The full description of each task is presented in Section 5. The $i^{\textrm{th}}$ multiplexed optical channel is detuned by $\Delta\omega_{i}$ from the angular frequency $\omega_{r_{i}}$ of its respective resonance. The frequency separation between the resonances corresponds to the free spectral range (FSR) of the MRR. The FSR can be approximately defined in terms of the wavelength ($\lambda$) as:
@@
where $R$ is the radius of the MRR (for the entirety of this work R = 7.5 $\mu$m). $n_{g}$ is the group index (no chromatic dispersion is considered). Throughout this work, if not specified, the frequencies of the exploited resonances for a number $M$ of optical channels are determined as follows:
@@ 4.2 Input layer
The input symbol sequence of each task, $u_{i}(n)$, is modulated at 1 GBd. $u_{i}(n)$ is masked by multiplying it with the sequence $m_{i}(n)$ to assign the random fixed weights of the RC. In TDRC systems, the masking sequence has a length equal to the desired number of virtual nodes, in this case, $N=50$ virtual nodes. Next, an optimized bias $\beta_{i}$ is added to the masked sequence as follows, resulting in $X_{i}(n)$:
@@
Each $u_{i}(n)$ and $m_{i}(n)$ are generated differently based on the mathematical definitions of each task, as further detailed in Section 5. For a 1 GBd input sequence, every virtual node has a duration of $\theta=\frac{1.0\textrm{ns}}{N}=20$ ps. The added biases, $\beta_{i}$, are necessary to fulfill one of the limits of the TCMT model used in this work as a small modulation index is assumed. The value of the bias for each task constrains the masked input signal to have a modulation index of less than 2%. Each optical channel has an average input power $\overline{P}_{\textrm{i}}$. The total power employed at the input of the MRR is then:
@@
The optical channels are then modulated by their respective masked input data sequence. The modulator (MOD) block in Fig. 1 refers to the encoding of the information in the intensity of the optical signal when pre-processing the data. Each optical channel is modulated in power linearly. Subsequently, the modulated optical channels are wavelength-multiplexed and injected into the input port of the MRR. The resulting electric field of the $i^{\textrm{th}}$ modulated channel is then defined as:
@@ 4.3 Reservoir layer
We consider a silicon-on-insulator (SOI) MRR with an input signal wavelength centered around 1550 nm. Given the bandgap of silicon, the electromagnetic field in such a wavelength range triggers the generation of free carriers via two-photon absorption (TPA) in the MRR cavity. The surge of excess carriers yields free carrier dispersion (FCD) which decreases the effective refractive index of the silicon, and results in a blue shift of the MRR resonance. Free-carrier absorption losses and the subsequent conversion of optical power to heat increase the temperature in the cavity. The resulting thermo-optic (TO) effect increases the refractive index and causes a red shift of the resonance. The nonlinear dynamics of FCD and the TO effect are a source of multistability and self-pulsing in the MRR, which have been extensively studied over the years (Johnson et al., 2006; Van Vaerenbergh et al., 2012; Zhang et al., 2013; Borghi et al., 2021).
@@
where the first term accounts for the loss rate due to the cavity waveguide attenuation ($\alpha$), followed by the loss rate due to the coupling with the two bus waveguides ($2/\tau_{c}$). The terms $\gamma_{\textrm{TPA/FCA}}$ refer to the loss rate due to FCA and TPA. Lastly, we define expressions for ($E_{\textrm{add}_{i}}$) and ($E_{\textrm{drop}_{i}}$):
@@
where $E_{\textrm{in}_{i}}$ is the $i^{\textrm{th}}$ electric field at the input port, and $\tau_{\textrm{d}}$ is the time delay added by the external feedback. Such time delay is assumed constant in the frequency range investigated, implicitly neglecting the impact of dispersion in the waveguide. $\kappa_{\textrm{d}}=0.95$ is the coupling factor of the delay waveguide. $\phi_{i}$ is the total phase shift experienced by each optical signal when propagated through the external waveguide, including the previously mentioned phase control $\Delta\phi$. It is defined as:
@@
The time constants $\tau_{\textrm{FC}}$ and $\tau_{\textrm{th}}$ are the lifetime of the carriers and the heat diffusion time constant, respectively. $\hbar$ is the reduced Planck’s constant and $m$ is the mass of the MRR. $\Gamma_{\textrm{FCA/th}}$ refer to the FCA and thermal confinement factors. $V_{\textrm{FCA/TPA}}$ denote the FCA and TPA effective volumes in the cavity. $\beta_{\textrm{TPA}}$ and $c_{\textrm{p}}$, are the silicon’s TPA coefficient, and specific heat, respectively. $\sigma_{\textrm{FCA}}$ is the cross-section of FCA in the cavity. The time-dependent term $P_{\textrm{abs}_{i}}(t)$ represents the total power absorbed in the MRR cavity and it is defined as follows:
@@
The system of differential equations in Eqs. 6, 12 and 13 is solved, by first normalizing the equations to dimensionless parameters and then using a conventional Runge-Kutta method. The discretization process of the signals into the steps of the Runge-Kutta solver follows the same procedure as done in (Giron Castro et al., 2024b). In this case, the mathematical procedure is just extended to account for the multiple electric fields.
The values of the simulated optical parameters are the same as in our initial study of this system (Giron Castro et al., 2024c) and are listed in Table 1.
@@ 4.4 Output layer
At the output of the reservoir (drop port of the MRR), the $M$ WDM signals are wavelength-demultiplexed and detected by individual photodiodes. The reservoir output for each optical channel $i$ and time-multiplexed node $j$ is determined at the photo-detection stage by calculating the modulus square of $E_{\textrm{drop}_{i,j}}(n)$.
@@
This brings additional nonlinearity through a square power function. Based on each detected signal, we train the output layer using ridge regression to calculate the $N$-size weight vector $W$ of each task. For the training of every multiplexed RC, we use a regularization parameter $\Lambda$ = $0.5\times 10^{-9}$, which was fine-tuned to obtain the best performance of the tasks for a given simulation in Section 7. The predicted sequence of each task $\hat{y_{i}}$(n) is consequently calculated as:
@@ 5 Benchmark tasks and computing metrics
In this section, we describe each of the benchmark tasks that have been used in this work: NARMA10 time-series prediction, signal waveform classification (SWC) as a basic classification task, and the equalization of a wireless channel affected by nonlinear distortions and noise, referred to as ChEq. The fourth task is concerned with the prediction of the future position of the target in an experimentally measured radar signal (IPIX). The details of the tasks were obtained from (Paquot et al., 2012; Duport et al., 2012).
@@ 5.1 NARMA-10 time-series prediction
As it is highlighted by Eq. 17, NARMA-10 is useful to test the memory capabilities of an RC scheme. For this task, $u_{0}(n)$ and $m_{0}(n)$ are uniformly distributed over the intervals [0.0, 0.5] and [0.0, +1.0], respectively. The metric performance is the normalized mean square error (NMSE) between the predicted and the target data sequences. The NMSE is defined as:
@@ 5.2 Signal waveform classification
In this toy classification task, the RC target is to differentiate correctly the shape of a waveform between square and sine waveforms as follows:
@@ 5.3 Wireless channel equalization
This task emulates a wireless channel model that is disturbed by multipath fading, noise, and high-order nonlinear distortions. The system is first modeled as a linear wireless channel with the input $d(n)$ undergoing the effect of multipath fading as shown in the following expression:
@@
This is followed by a combination of second-order and third-order nonlinear distortion besides the addition of pseudo-random additive Gaussian noise with zero mean, denoted by $v(n)$. The resulting distorted sequence becomes the input of the RC:
@@
The target of the RC is to reconstruct the original signal, $d(n)$, which is an independent, identically distributed random sequence with values taken from $\{3,-1,+1,+3\}$.
During the pre-processing stage, we added a bias to the input, $u(n)+5$, before masking the signal. The values of the masking sequence, $m_{2}(n)$ are taken from a uniform distribution in the interval $[-1,+1]$. During the training, the target sequence is shifted in time by 2 before minimizing the square error $(\hat{y}(n)-d[n-2])^{2}$. From the obtained output, the values of the predicted sequence $\hat{y}_{2}(n)$ are approximated to their nearest neighbor from the values $\{3,-1,+1,+3\}$. The performance metric is the symbol error ratio (SER) between the original and predicted sequences. We use a signal-to-noise ratio (SNR) of 32 dB.
@@ 5.4 IPIX radar task
This task consists of a $K$-step time-series prediction of a signal. The target is an experimental backscattered radar signal from the ocean surface, which was measured by the McMaster University IPIX radar. Among the available datasets (Haykin, 2001) we used the ones related to the average height of waves. More specifically, the one corresponding to the high sea states (average wave height of 1.8 meters). The dataset is generated from the in-phase and quadrature outputs of the radar IQ demodulator. The performance is determined by calculating the NMSE between the prediction and the experimental measurements of the signal. We simulated the system for $K=2$. The 2-D signal is flattened and then processed sequentially by the RC. The training and testing sets used in this task are arbitrarily selected from the original datasets.
@@ 5.5 Linear, nonlinear and total memory capacity
RC is capable of buffering a finite number of previous inputs. How long an input drives or has an influence on the reservoir state is limited, which reduces the computing load in RC schemes. However, this also limits the number of past inputs that it can accurately predict. Therefore, we can quantify the memory capacity ($MC$) of RC by evaluating the ability of a particular RC scheme to reconstruct its own input $K$ steps in the past. As first defined by (Jaeger, 2002), the linear $MC$ of RC can be calculated by training it to reconstruct the input (usually uniformly distributed) given by $y_{K}(n)=u(n-K)$. The mathematical expression for a specific value of $K$ is defined as:
@@
There is a limit to the significant $K$ steps contributing to the memory capacity of a given RC scheme. In (Dambre et al., 2012) it is demonstrated to be equal to the number of nodes $N$, so that the value used for $K_{\textrm{max}}$ in this work is 50. The total linear $MC$ is defined as the sum of the capacities considering all the values of $K\leq K_{\textrm{max}}$:
@@
The previous definition accounts only for the linear $MC$ of the system. In order to expand the definition to higher-order functions we can generalize the concept as the reconstruction of a set of basis functions in the Hilbert space of fading memory functions (Dambre et al., 2012; Hülser et al., 2023). The nonlinear memory capacity is calculated by using a basis out of finite products of higher-order Legendre polynomials. The basis function of the $i^{\textrm{th}}$-order is noted as $\mathbb{P}_{i}$, for $y_{i}(n)=\mathbb{P}_{i}$. Thus, the nonlinear capacities are defined as follows:
@@
Then, the total memory capacity $MC$, or the so-called, information processing capacity, (Jaeger, 2002; Hülser et al., 2023) is calculated by summing-up all capacities of the $i^{\textrm{th}}$ order up to the highest order of function considered, $H$:

### v2

@@ 3 Introduction
Over recent years, the development of photonic computing has attracted significant interest within the scientific community and positioned photonics as a potential technology for novel computing schemes (Huang et al., 2022; McMahon, 2023). Light possesses unique properties that can be harnessed for computing such as the capability for massive parallelization. This can be achieved by multiplexing several optical channels using different wavelengths within an available range of the electromagnetic spectrum. This is a well-established technique used for optical communication known as wavelength division multiplexing (WDM). In WDM, a high density of channels can be transmitted through the same medium with relatively low interference between them, ultimately boosting the total communication capacity as several data streams can be transmitted simultaneously. If a similar principle is used for computing, this can be translated into different computing tasks being addressed simultaneously.
@@
These studies on MRR-based TDRC however, focused on investigating the system when addressing a single task at a time. This is where the benefits of WDM can come into play to enhance the computing capacity of photonic RC schemes. For instance, in (Gooskens et al., 2022), the wavelength dimension is used in a photonic RC scheme to enhance the performance of a specific task (nonlinear signal equalization and boolean operations). Similarly, in (Li et al., 2023) WDM is used in an RC based on a Fabry-Perot semiconductor laser to achieve superior throughput and performance in the task of signal equalization in an optical fiber communication link. Lastly, in (Lupo et al., 2023) a photonic RC based on phase modulation and spectral filtering demonstrated the simultaneous operation of two different instances of the same task, which are encoded using frequency combs. In an initial study (Giron Castro et al., 2024), we combined WDM with MRR-based TDRC to show that this scheme also offers the potential of parallel computing by simultaneously addressing three distinct computing tasks.
In this work, we extend the study in (Giron Castro et al., 2024) by systematically analyzing the memory and nonlinear capabilities of the proposed WDM MRR-based TDRC scheme and investigating the impact of additional computing tasks in parallel. Finally, we quantify the effect of the phase control in the external waveguide that is used in this scheme as a feedback mechanism. The numerical model of the system is based on the well-studied temporal coupled-mode theory (TCMT).
@@ 4 Multitask WDM MRR-based TDRC
On a high level, the TDRC scheme based on an MRR relies on injecting the input data encoded onto an optical carrier into the MRR at its input port and detecting a non-linearly transformed version at the drop port. An additional feedback connection with an optimized delay is normally included between the through and add port of the MRR to provide additional memory (Donati et al., 2022, 2024; Giron Castro et al., 2024b).
@@ 4.1 Setup and frequency allocation
The simulated system is shown in Fig. 1. We consider $M$ wavelength-multiplexed optical channels, each modulated by the masked input sequence corresponding to a particular task. The full description of each task is presented in Section 5. The $i^{\textrm{th}}$ multiplexed optical channel is detuned by $\Delta\omega_{i}$ from the angular frequency $\omega_{r_{i}}$ of its respective resonance. The frequency separation between the resonances corresponds to the free spectral range (FSR) of the MRR. The FSR can be approximately defined in terms of the wavelength ($\lambda$) as expressed in Eq. 1:
@@
where $R$ is the radius of the MRR (for the entirety of this work R = 7.5 $\mu$m). $n_{g}$ is the group index (no chromatic dispersion is considered). Throughout this work, if not specified, the frequencies of the exploited resonances for a number $M$ of optical channels are determined as follows expressed in Eq. 2:
@@ 4.2 Input layer
The input symbol sequence of each task, $u_{i}(n)$, is modulated at 1 GBd. $u_{i}(n)$ is masked by multiplying it with the sequence $m_{i}(n)$ to assign the random fixed weights of the RC. In TDRC systems, the masking sequence has a length equal to the desired number of virtual nodes, in this case, $N=50$ virtual nodes. Next, an optimized bias $\beta_{i}$ is added to the masked sequence as follows in Eq. 3, resulting in $X_{i}(n)$:
@@
Each $u_{i}(n)$ and $m_{i}(n)$ are generated differently based on the mathematical definitions of each task, as further detailed in Section 5. For a 1 GBd input sequence, every virtual node has a duration of $\theta=\frac{1.0\textrm{ns}}{N}=20$ ps. The added biases, $\beta_{i}$, are necessary to fulfill one of the limits of the TCMT model used in this work as a small modulation index is assumed. The value of the bias for each task constrains the masked input signal to have a modulation index of less than 2%. Each optical channel has an average input power $\overline{P}_{\textrm{i}}$. The total power employed at the input $(\overline{P}_{\textrm{T}}$, as expressed in Eq. 4)))) of the MRR is then:
@@
The optical channels are then modulated by their respective masked input data sequence. The modulator (MOD) block in Fig. 1 refers to the encoding of the information in the intensity of the optical signal when pre-processing the data. Each optical channel is modulated in power linearly. Subsequently, the modulated optical channels are wavelength-multiplexed and injected into the input port of the MRR. The resulting electric field of the $i^{\textrm{th}}$ modulated channel is then defined as in Eq. 5:
@@ 4.3 Reservoir layer
We consider a silicon-on-insulator (SOI) MRR with an input signal wavelength centered around 1550 nm. Given the bandgap of silicon, the electromagnetic field in such a wavelength range triggers the generation of free carriers via two-photon absorption (TPA) in the MRR cavity. The surge of excess carriers yields free carrier dispersion (FCD), which decreases the effective refractive index of the silicon, and results in a blue shift of the MRR resonance. Free-carrier absorption losses and the subsequent conversion of optical power to heat increase the temperature in the cavity. The resulting thermo-optic (TO) effect increases the refractive index and causes a red shift of the resonance. The nonlinear dynamics of FCD and the TO effect are a source of multistability and self-pulsing in the MRR, which have been extensively studied over the years (Johnson et al., 2006; Van Vaerenbergh et al., 2012; Zhang et al., 2013; Borghi et al., 2021).
@@
where the first term accounts for the loss rate due to the cavity waveguide attenuation ($\alpha$), followed by the loss rate due to the coupling with the two bus waveguides ($2/\tau_{c}$). The terms $\gamma_{\textrm{TPA/FCA}}$ refer to the loss rate due to FCA and TPA. Lastly, in Eqs. 9 and 10 we define expressions for ($E_{\textrm{add}_{i}}$) and ($E_{\textrm{drop}_{i}}$):
@@
where $E_{\textrm{in}_{i}}$ is the $i^{\textrm{th}}$ electric field at the input port, and $\tau_{\textrm{d}}$ is the time delay added by the external feedback. Such time delay is assumed constant in the frequency range investigated, implicitly neglecting the impact of dispersion in the waveguide. $\kappa_{\textrm{d}}=0.95$ is the coupling factor of the delay waveguide. $\phi_{i}$ is the total phase shift experienced by each optical signal when propagated through the external waveguide, including the previously mentioned phase control $\Delta\phi$. It is defined in Eq. 11 as:
@@
The time constants $\tau_{\textrm{FC}}$ and $\tau_{\textrm{th}}$ are the lifetime of the carriers and the heat diffusion time constant, respectively. $\hbar$ is the reduced Planck’s constant and $m$ is the mass of the MRR. $\Gamma_{\textrm{FCA/th}}$ refer to the FCA and thermal confinement factors. $V_{\textrm{FCA/TPA}}$ denote the FCA and TPA effective volumes in the cavity. $\beta_{\textrm{TPA}}$ and $c_{\textrm{p}}$, are the silicon’s TPA coefficient, and specific heat, respectively. $\sigma_{\textrm{FCA}}$ is the cross-section of FCA in the cavity. The time-dependent term $P_{\textrm{abs}_{i}}(t)$ represents the total power absorbed in the MRR cavity and it is defined by Eq. 14:
@@
The system of differential equations in Eqs. 6, 12 and 13 is solved, by first normalizing the equations to dimensionless parameters and then using a conventional Runge-Kutta method. The discretization process of the signals into the steps of the Runge-Kutta solver follows the same procedure as done in (Giron Castro et al., 2024b). In this case, the mathematical procedure is just extended to account for the multiple electric fields. The values of the simulated optical parameters are the same as in our initial study of this system (Giron Castro et al., 2024) and are listed in Table 1. We calculate the values of the quality factor ($Q$) and the full width at half-maximum (FWHM) of the silicon MRR, as defined in (Biasi et al., 2024).
@@ 4.4 Output layer
At the output of the reservoir (drop port of the MRR), the $M$ WDM signals are wavelength-demultiplexed and detected by individual photodiodes. The reservoir output for each optical channel $i$ and time-multiplexed node $j$ is determined at the photo-detection stage by calculating the modulus square of $E_{\textrm{drop}_{i,j}}(n)$ as in Eq. 15.
@@
This brings additional nonlinearity through a square power function. Based on each detected signal, we train the output layer using ridge regression to calculate the $N$-size weight vector $W$ of each task. For the training of every multiplexed RC, we use a regularization parameter $\Lambda$ = $1.0\times 10^{-9}$, which was fine-tuned to obtain the best performance of the tasks for a given simulation in Section 7. The predicted sequence of each task $\hat{y_{i}}$(n) is consequently calculated as in Eq. 16:
@@ 5 Benchmark tasks and computing metrics
In this section, we describe each of the benchmark tasks that have been used in this work: NARMA-10 time-series prediction, signal waveform classification (SWC) as a basic classification task, and the equalization of a wireless channel affected by nonlinear distortions and noise, referred to as ChEq. The fourth task is concerned with the prediction of the future position of the target in an experimentally measured radar signal (IPIX). The details of the tasks were obtained from (Paquot et al., 2012; Duport et al., 2012).
@@ 5.1 NARMA-10 time-series prediction
As it is highlighted by Eq. 17, NARMA-10 is useful to test the memory capabilities of an RC scheme. For this task, $u_{0}(n)$ and $m_{0}(n)$ are uniformly distributed over the intervals [0.0, 0.5] and [0.0, +1.0], respectively. The metric performance is the normalized mean square error (NMSE) between the predicted and the target data sequences. The mathematical definition of the NMSE is given by Eq. 18:
@@ 5.2 Signal waveform classification
In this toy classification task, the RC target is to differentiate correctly the shape of a waveform between square and sine waveforms as determined by Eq. 19:
@@ 5.3 Wireless channel equalization
This task emulates a wireless channel model that is disturbed by multipath fading, noise, and high-order nonlinear distortions. The system is first modeled as a linear wireless channel with the input $d(n)$ undergoing the effect of multipath fading as shown in Eq. 20:
@@
This is followed by a combination of second-order and third-order nonlinear distortion besides the addition of pseudo-random additive Gaussian noise with zero mean, denoted by $v(n)$. The resulting distorted sequence becomes the input of the RC, mathematically expressed in Eq. 21:
@@
The target of the RC is to reconstruct the original signal, $d(n)$, which is an independent, identically distributed random sequence with values taken from $\{-3,-1,+1,+3\}$.
During the pre-processing stage, we added a bias to the input, $u(n)+20$, before masking the signal. The values of the masking sequence, $m_{2}(n)$ are taken from a uniform distribution in the interval $[-1,+1]$. During the training, the target sequence is shifted in time by 2 before minimizing the square error $(\hat{y}(n)-d[n-2])^{2}$. From the obtained output, the values of the predicted sequence $\hat{y}_{2}(n)$ are approximated to their nearest neighbor from the values $\{-3,-1,+1,+3\}$. The performance metric is the symbol error ratio (SER) between the original and predicted sequences. We use a signal-to-noise ratio (SNR) of 32 dB.
@@ 5.4 IPIX radar task
This task consists of a $K$-step time-series prediction of a signal. The target is an experimental backscattered radar signal from the ocean surface, which was measured by the McMaster University IPIX radar. Among the available datasets (Haykin, 2001) we used the ones related to the average height of waves. More specifically, the one corresponding to the high sea states (average wave height of 1.8 meters). The dataset is generated from the in-phase and quadrature outputs of the radar IQ demodulator. The performance is determined by calculating the NMSE between the prediction and the experimental measurements of the signal. We simulated the system for $K=2$. The 2-D signal is flattened and then processed sequentially by the RC. The training and testing sets used in this task were arbitrarily selected from the original datasets.
@@ 5.5 Linear, nonlinear and total memory capacity
RC is capable of buffering a finite number of previous inputs. How long an input drives or has an influence on the reservoir state is limited, which reduces the computing load in RC schemes. However, this also limits the number of past inputs that it can accurately predict. Therefore, we can quantify the memory capacity ($MC$) of RC by evaluating the ability of a particular RC scheme to reconstruct its own input $K$ steps in the past. As first defined by (Jaeger, 2002), the linear $MC$ of RC can be calculated by training it to reconstruct the input (usually uniformly distributed) given by $y_{K}(n)=u(n-K)$. The mathematical expression for a specific value of $K$ is defined by Eq. 22:
@@
There is a limit to the significant $K$ steps contributing to the memory capacity of a given RC scheme. In (Dambre et al., 2012) it is demonstrated to be equal to the number of nodes $N$, so that the value used for $K_{\textrm{max}}$ in this work is 50. The total linear $MC$ is defined as the sum of the capacities considering all the values of $K\leq K_{\textrm{max}}$ as expressed by Eq. 23:
@@
The previous definition accounts only for the linear $MC$ of the system. In order to expand the definition to higher-order functions we can generalize the concept as the reconstruction of a set of basis functions in the Hilbert space of fading memory functions (Dambre et al., 2012; Hülser et al., 2023). The nonlinear memory capacity is calculated by using a basis out of finite products of higher-order Legendre polynomials. The basis function of the $i^{\textrm{th}}$-order is noted as $\mathbb{P}_{i}$, for $y_{i}(n)=\mathbb{P}_{i}$. Thus, the nonlinear capacities are defined by Eq. 24 as follows:
@@
Then, the total $MC$, or the so-called, information processing capacity, (Jaeger, 2002; Hülser et al., 2023) is calculated by summing-up all capacities of the $i^{\textrm{th}}$ order up to the highest order of function considered, $H$ (Eq. 25):